Please, I need some help with these problems. If anyone can help please.
1. {4x+y=13
{2x-y=5
2. {2x+4y=6
{5x+10=15
3. {x+y-z=-1
{-4x-y+2z=-7
{2x-2y-5z=7

Answer by stanbon(75887)

1. {4x+y=13
{2x-y=5
--------
Solving with substitution:
Solve the 2nd equation as follows:
y = 2x-5
Substitute into the 1st equation to solve for x:
4x+(2x-5) = 13
6x=18
x=3
------------
Substitue into y=2x-5 to solve for y:
y=2(3)-5
y = 1
---------
Check the answers in 4x+y=13
4*3 + 1 = 13
13 = 13
------------------------------------------------

2. {2x+4y=6
{5x+10=15
---------------
Multiply 1st by 5 and 2nd by 2 to get:
10x+20y=30
10x+20y=30
-----------------
This shows that each of the equation is a multiple of the other.
The solution is all points on the line where
y = (-1/2)x+(3/2)
-----------------------
3. {x+y-z=-1
{-4x-y+2z=-7
{2x-2y-5z=7
These are too messy to work online but using matrix methods
I can tell you the solution is:
x=3 ;y = -3 ; z = 1
================
Cheers,
Stan H.